OpenStudy (anonymous):

Solving for x... Can anybody help? I think I did it right but want to make sure. Log2 base 2 ^ x-5 = 5

OpenStudy (anonymous):

Do you mean\[\log_22^{x-5}\]?

OpenStudy (anonymous):

If so, that's\[\log_22^{x-5}=(x-5)\log_22=(x-5).1=x-5\]

OpenStudy (anonymous):

Yes that is the equation. Solving for X.

OpenStudy (anonymous):

what is X?

OpenStudy (anonymous):

Well, you can't solve for x unless \[\log_22^{x-5}=(something)\]

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

If there's no '=' sign in your question, it's an 'expression' that you need to simplify.

OpenStudy (anonymous):

x=10 does that help?

OpenStudy (anonymous):

Sorry - I didn't see the 5

OpenStudy (anonymous):

\[x-5=5 \rightarrow x=10\]

OpenStudy (anonymous):

Yes It does . Thank you . That is what I got.